# El Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST data

### v1

@@ 4.7 Model families
The third “Cluster+Shear” model is a combination of the two models above as it contains both an external shear component and the cluster’s deflection field. We consider this more advanced model as our fiducial one. The external shear in this case is expected to capture additional azimuthal structures that have not been included in the cluster lens model, such as nearby faint galaxies that are visible in some NIRCam filters. However, as recently recalled by Etherington et al. (2023b), there is no guarantee that external shear captures only structures external to the lens, but can instead compensate for the lack of azimuthal freedom in the main deflectors mass model. In our case, one can also interpret the external shear as a local correction to the cluster deflection field.
@@ 5.1 Overall fit to the multi-band NIRCam data
Figure 7: Mean lens model among the best-fit posterior samples for our fiducial “Cluster+Shear” model. First row, from left to right: color image of the data, color image of the model with reduced $\chi^{2}$ from image residuals (which differs from the full log-likelihood, see Eq. 4.5), convergence map (with convergence and surface brightness contours), and color image of the unlensed source model. Remaining rows, from left to right: data in the indicated filter, model, normalized residuals (areas excluded from the fit appear white), unlensed source model, S/N source map defined as the mean source model divided by the standard deviation of the model posterior in each source pixel. Critical lines and caustics are indicated in some of the panels as white solid line in the image model panel, and dotted lines in the right column panels.
@@
In the image plane, the model fits remarkably well the data in all bands. The lensed arcs, although containing a large number of features at difference scales, are particularly well fitted. We also note the dynamic range of the source model, which spans more than three orders of magnitudes from fainter structures in filter F150W to the brighter central region in filter F [^4] W. The surface brightness of the L2 and L3 galaxies are also accurately modeled. Small scale residuals that still remain close to L2 and L3 centroids are likely caused by slightly inaccuracies in the PSF model, as they are similar to those of quasar-host separation analyses (e.g., see Fig. 2 of Ding et al. (2022) or Fig. 4 of Stone et al. (2023)).
@@
We distinguish two main categories of residuals. First, the surface brightness of L1 is not accurately reproduced by the model close to the center of the galaxy and beyond the arc in the norther part of the cutout (see Appendix A.1). Second, the brightest part of the southern arc (corresponding to the very center of the source galaxy), as well as the compact feature (almost point-like) at the north of L2, are not fully captured by the source model. While the southern part leads to mild residuals only in F [^4] W (below the $3\sigma$ level), the compact feature leads to strong residuals (more than $6\sigma$ within some pixels) with some symmetric features. In the source plane, this clump is clearly offset and distinct from the center of the source galaxy, as shown in the rightmost panel of Fig. 1. We believe this clump is resolved (as opposed to a point source), because increasing the source model resolution allows us to capture its extent over several pixels. We discuss in more detail this specific feature in Sect. 4.
@@
The multi-band model of the background galaxy reveals a heavily wavelength-dependent morphology. In F150W and F200W filters, most of the detected regions of the source are located outside the main astroid (a.k.a., tangential) caustic. Interestingly, the morphology in these filters do not reveal any significantly bright component that could be associated to a centroid. It is only in F277W and redder bands that we see a more symmetric morphology containing features that could be attributed to a brighter central bulge, dust attenuated areas, or regions with lower stellar density. Moreover, our model reveals new features located south of the bulge, which may be associated to spiral arms. The source S/N is, as expected, higher along the caustics, because these regions are highly magnified and have at least two multiple images visible in the data. The bulge of the source galaxy is the highest S/N region and is located remarkably close to the southern part of the main tangential caustic (which features a double-cusp shape). Coincidentally, the upper part of the source crosses the northern cusp of the same caustic. Overall, the source S/N is larger inside and along the caustics, but is not significantly lower outside (as one may expect). We attribute this globally uniform S/N to the significant brightness outside the caustics in the mid-infrared, which effectively leads to higher S/N at these locations.
@@
Our source S/N maps also allow us to discard highly uncertain source plane features, such as the smooth blob located around $(-27\aas@@fstack{\prime\prime}8,-5\aas@@fstack{\prime\prime}6)$. These spurious features are in reality caused by the leakage of un-modeled lens light features that contaminate the source model. Indeed, the purely smooth light model of L1 is inaccurate in some locations that overlap with the arc mask, causing the lensed source model to predict non-zero flux along the edge source plane grid. However, the posterior uncertainty of the source pixels is large, as expected. We expect that a better model of the deflectors surface brightness will remove these low-significance features from the source model. While a more flexible deflectors light model may be warranted in the future, our current model does not support the hypothesis of possible over- or under-subtraction of deflectors light over the arc mask region (i.e., all additional amplitudes varied over the arc mask region in each band remain extremely low $\lesssim 10^{5}$ ${}^{-}$/s).
@@
While designing the arc mask, we noticed one luminous structure on the West side of the arc around image plane position $(2\aas@@fstack{\prime\prime}8,-0\aas@@fstack{\prime\prime}7)$ appearing bluer than the rest of the arc in the NIRCam data. This color difference may be the sign that this structure does not belong to the source, but is located along the line-of-sight instead. However, due to its close proximity with the arc and the absence of objective evidence that it does not belong to the source, we did not mask it and thus allow our source model to capture it. The unlensed version of this structure is clearly visible in the fourth column of Fig. 7 as a thin elongated feature of size $\sim 0\aas@@fstack{\prime\prime}1$ around position $(-26\aas@@fstack{\prime\prime}4,-5\aas@@fstack{\prime\prime}2)$. We note that Kamieneski et al. (2023) also did not mask this structure and ray-traced it to their source plane model (visible on the top right panel of their Fig. 2). Moreover, this structure being faint in the data, it has likely a small mass which may explain why we do not see signs of local distortions in the arc.
@@ 5.2 Critical lines and caustics
Figure 8: Critical curves and caustics predicted by the different lens models explored in this work, namely “Shear only”, “Cluster only” and our fiducial model “Cluster+Shear” that best fits the data. Also shown are the critical lines and caustics from recent models of El Anzuelo, namely from Kamieneski et al. (2023, constrained by pairs of multiple images with JWST imaging data) and Caminha et al. (2023, cluster-scale model of El Gordo using HST and MUSE data sets). Caustics from our models are shown with dashed lines for clarity. The left panel is focused on the image plane, while the right panel is focused on the source region predicted by models including the cluster mass distribution. Note how the position of the caustics for the “Shear only” model is aligned with the lens (left panel), compared to the caustics of other models (right panel). The predicted (unlensed) surface brightness of the lens and source galaxies from the best-fit “Cluster+Shear” model (red curves) is shown in the background, and North and East directions are indicated by the arrows.
@@
Although using pairs of multiply imaged features as constraints, the lens model of Kamieneski et al. (2023) is qualitatively very similar to ours. We show their critical lines and caustics [^11] An offset of $+0\aas@@fstack{\prime\prime}25\ {\rm E},+0\aas@@fstack{\prime\prime}01\ {\rm N}$ have been added to the critical lines and caustics of Kamieneski et al. (2023) for proper comparison with our models (P. Kamieneski, private communication). We have also found a similar offset between deflectors positions reported in Caminha et al. (2023) and our fitted light profile positions. in comparison to ours in Fig. 8. As expected, the main differences in the critical lines are directly related to the amount and location of the multiple images used in Kamieneski et al. (2023). A striking example is the westernmost part of the arc—around position $(2\aas@@fstack{\prime\prime}2,1^{\prime\prime})$—, for which Kamieneski et al. (2023) used a pair of multiple images which are probably unresolved by JWST and appears as point sources (images 2b and 2c in Fig. 11). In between these two images, all critical lines are overlapping, showing that both the arc and point-like constraints locally give similar results, despite our extended source model not including any point source components. On the contrary, in locations where either pairs of images are missing (e.g., south-western part of the arc) or inaccurate multiple images are used (northern part of the arc), clear differences do arise. In particular, the zoom-in inset in Fig. 8 focuses on a region with many multiple images of central regions of the source galaxy and shows deviations between the critical lines from different models. As we discuss in more detail in Sect. 5.4, we find that one feature in the arc may have been incorrectly attributed to a lensed image of the bright source bulge. We believe that this incorrect attribution causes the deviation between critical lines in the top right corner of the inset in Fig. 11, despite a region with a large number of lensing constraints. Focusing now on the caustics, we see a clear offset in the main astroid caustic between the model of Kamieneski et al. (2023) and our “Shear only” model. This result may be surprising at first, since both models are similarly parametrized (power-law profiles embedded in an external shear). However, this offset could be explained by the fact that our models including external shear all predict a non-zero mass for L3, as well as differences in the lensing constraints between the two approaches.
@@
We also compare in Fig. 8 our critical curves and caustics to the ones predicted by the cluster model. The cluster model was constructed without any constraints from El Anzuelo hence L1, L2 and L3 are mainly constrained through scaling relations based on their photometry (Caminha et al. 2023). We find that critical lines between our galaxy-scale model and this cluster-scale overall agree well, in particular in the vicinity L1 and L2. As shown in the inset of Fig. 8, the characteristic feature in the critical lines is comparable, meaning that qualitatively, the relative masses of L1 and L2 are consistent among the models. The cluster model displays rounder critical lines than our model, as expected from the use of spherically symmetric mass profiles. The discrepancy is the largest in the vicinity of L3, as this galaxy lies further away from the arcs. The cluster mass model assigns a larger mass to L3, which is also visible in source plane, where the eastern astroid caustic is significantly larger than in our models.
@@ 5.3 Constraints on mass model parameters
121212 The above constraints are computed from the marginalized posterior distributions of our fiducial “Cluster+Shear” model.
@@
\$$\star$\$\$\$$\star$\$\$footnotetext: Except for $\theta_{\rm E}$, all L3 parameters are fixed to an isothermal slope ($\gamma=2$) or to values informed by the surface brightness model.
In this section we present our quantitative constraints on the mass distribution of the main deflectors. We show in Fig. 9 the joint posterior distributions of a subset of key mass model parameters, for each of the three model families (see Appendix A.2 for the full corner plot). We also summarize in Table 2 all constraints on mass model parameters for the fiducial “Cluster+shear” model. As expected, the Einstein radius of L1 and L2 as well the external shear strength significantly decrease after including the cluster mass distribution. Unlike the model of Kamieneski et al. (2023) composed of SIE profiles embedded in an external shear (which is the most similar to our “Shear only” model) the mass of L3 is constrained by the imaging data, except for the “Cluster only” model family. Our fiducial model predicts $\theta_{\rm E,L3}=0\aas@@fstack{\prime\prime}32\pm 0\aas@@fstack{\prime\prime}03$, which is $\sim 3\sigma$ away from the stellar kinematics estimate $\theta_{\rm E,SIS}=0\aas@@fstack{\prime\prime}49\pm 0\aas@@fstack{\prime\prime}05$. Therefore, the prior from stellar kinematics appears rather unconstraining compared to the imaging data, despite the angular distance between L3 and the lensing features.
@@
In Appendix A.3 we extensively check the robustness of our constraints against different modeling choices. These checks include the choice of VI sampler, the effect of inaccurate lens light model and the uncertainty associated to the cluster-scale model. In particular, we note that our models very weakly dependent on the source resolution, both in terms of constraints on mass model parameters and morphology of the reconstructed source. We believe this low dependence on the source resolution is a direct consequence of the efficient regularization of our correlated field source model, which allows us to reach high-resolution without compromising on computation time and model quality.
Finally, we find our lens model to be sensitive to changes in the cluster-scale mass distribution, as drawing different posterior samples from the cluster model of Caminha et al. (2023) lead to measurable changes in $\chi^{2}_{\nu}$. In particular, some posterior samples from the cluster model lead to a lower $\chi^{2}_{\nu}$ than with the best-fit cluster model we use in our fiducial model. Therefore, the El Anzuelo lensing features may be used to discriminate between different cluster models, thus providing additional constraints on the mass distribution of large-scale components of El Gordo (i.e., its dark matter halo). Although this particular result demands further investigation, it shows that extended arcs, although primarily constraining the structure of individual cluster members, may also provide useful constraints on the mass distribution at the scale of the cluster.

### v2

@@ 4.7 Model families
The third “Cluster+Shear” model is a combination of the two models above as it contains both an external shear component and the cluster’s deflection field. We consider this more advanced model as our fiducial one. The external shear in this case is expected to capture additional azimuthal structures that have not been included in the cluster lens model, such as nearby faint galaxies that are visible in some NIRCam filters. However, as recently recalled by Etherington et al. (2024), there is no guarantee that external shear captures only structures external to the lens, but can instead compensate for the lack of azimuthal freedom in the main deflectors mass model. In our case, one can also interpret the external shear as a local correction to the cluster deflection field.
@@ 5.1 Overall fit to the multi-band NIRCam data
Figure 7: Mean lens model among the best-fit posterior samples for our fiducial “Cluster+Shear” model. First row, from left to right: color image of the data, color image of the model with reduced $\chi^{2}$ from image residuals (which differs from the full log-likelihood, see Eq. 4.5), convergence map (with convergence and surface brightness contours), and color image of the unlensed source model. Remaining rows, from left to right: data in the indicated filter, model, normalized residuals (areas excluded from the fit appear white), unlensed source model, S/N source map defined as the mean source model divided by the standard deviation of the model posterior in each source pixel. Critical lines and caustics are indicated in some of the panels as white solid line in the image model panel, and dotted lines in the right column panels. The source S/N is higher within and along the caustics, since image multiplicity and lensing magnification are higher, respectively.
@@
In the image plane, the model fits remarkably well the data in all bands. The lensed arcs, although containing a large number of features at difference scales, are particularly well fitted. We also note the dynamic range of the source model, which spans more than three orders of magnitudes from fainter structures in filter F150W to the brighter central region in filter F [^4] W. The surface brightness of the L2 and L3 galaxies are also accurately modeled. Small scale residuals that still remain close to L2 and L3 centroids are likely caused by slight inaccuracies in the PSF model, as they are similar to those of quasar-host separation analyses (e.g., see Fig. 2 of Ding et al. (2022) or Fig. 4 of Stone et al. (2023)).
@@
We distinguish two main categories of residuals. First, the surface brightness of L1 is not accurately reproduced by the model close to the center of the galaxy and beyond the arc in the norther part of the cutout (see Appendix C.1). Second, the brightest part of the southern arc (corresponding to the very center of the source galaxy), as well as the compact feature (almost point-like) at the north of L2, are not fully captured by the source model. While the southern part leads to mild residuals only in F [^4] W (below the $3\sigma$ level), the compact feature leads to strong residuals (more than $6\sigma$ within some pixels) with some symmetric features. In the source plane, this clump is clearly offset and distinct from the center of the source galaxy, as shown in the rightmost panel of Fig. 1. We believe this clump is resolved (as opposed to a point source), because increasing the source model resolution allows us to capture its extent over several pixels. We discuss in more detail this specific feature in Sect. 5.5.
@@
The multi-band model of the background galaxy reveals a heavily wavelength-dependent morphology. In F150W and F200W filters, most of the detected regions of the source are located outside the main astroid (a.k.a., tangential) caustic. Interestingly, the morphology in these filters do not reveal any significantly bright component that could be associated to a centroid. It is only in F277W and redder bands that we see a more symmetric morphology containing features that could be attributed to a brighter central bulge, dust attenuated areas, or regions with lower stellar density. Moreover, our model reveals new features located south of the bulge, which may be associated to spiral arms.
The source S/N is, as expected, higher along the caustics, because these regions are highly magnified and have at least two multiple images visible in the data. The bulge of the source galaxy is the highest S/N region and is located remarkably close to the southern part of the main tangential caustic (which features a double-cusp shape). Coincidentally, the upper part of the source crosses the northern cusp of the same caustic. Overall, the source S/N is larger inside and along the caustics, but is not significantly lower outside (as one may expect). We attribute this globally uniform S/N to the significant brightness outside the caustics in the mid-infrared, which effectively leads to higher S/N at these locations. We also investigate if some features in the source S/N map can be a consequence of the spatially varying exposure time (see Fig. 6) and find that lower S/N regions (after being lensed) do not overlap with lower exposure time regions. Therefore, the source S/N is dominated by lensing magnification and the intrinsic source brightness, and the non-uniform exposure time does not impact the reconstructed source.
@@
Our source S/N maps also allow us to discard highly uncertain source plane features, such as the smooth blob located around $(-27\aas@@fstack{\prime\prime}8,-5\aas@@fstack{\prime\prime}6)$. These spurious features are in reality caused by the leakage of un-modeled lens light features that contaminate the source model. Indeed, the purely smooth light model of L1 is inaccurate in some locations that overlap with the arc mask, causing the lensed source model to predict non-zero flux along the edge source plane grid. However, the posterior uncertainty of the source pixels is large, as expected. We expect a better model of the deflectors surface brightness to remove these low-significance features from the source model. While a more flexible deflectors light model may be warranted in the future, our current model does not support the hypothesis of possible over- or under-subtraction of deflectors light over the arc mask region (i.e., all additional amplitudes varied over the arc mask region in each band remain extremely low $\lesssim 10^{5}$ e-/s).
@@
While designing the arc mask, we noticed one luminous structure on the West side of the arc around image plane position $(2\aas@@fstack{\prime\prime}8,-0\aas@@fstack{\prime\prime}7)$ appearing bluer than the rest of the arc in the NIRCam data. This color difference may be the sign that this structure does not belong to the source, but is located along the line-of-sight instead. We also notice that this bluer structure has similar colors as the two galaxies located West to the arc (visible in the color image of Fig. 1). Moreover, the foreground ($z=0.63$) galaxy group reported in Caminha et al. (2023) contains galaxies that appear bluer than El Gordo members. These blue objects in the vicinity of El Anzuelo may thus be low-mass or dwarf galaxies lying at a similar redshift as the foreground galaxy group. Unfortunately, these objects are not detected in the MUSE data, such that we are unable to confirm their nature. Therefore, due to its close proximity with the arc and the absence of objective evidence that it does not belong to the source, we did not mask it and thus allow our source model to capture it. The unlensed version of this structure is clearly visible in the fourth column of Fig. 7 as a thin elongated feature of size $\sim 0\aas@@fstack{\prime\prime}1$ around position $(-26\aas@@fstack{\prime\prime}4,-5\aas@@fstack{\prime\prime}2)$. We note that Kamieneski et al. (2023) also did not mask this structure and ray-traced it to their source plane model (visible on the top right panel of their Fig. 2). Moreover, this structure being faint in the data, it has likely a small mass which may explain why we do not see signs of local distortions in the arc.
@@ 5.2 Critical lines and caustics
Figure 8: Critical curves and caustics predicted by the different lens models explored in this work, namely “Shear only”, “Cluster only” and our fiducial model “Cluster+Shear” that best fits the data. Also shown are the critical lines and caustics from recent models of El Anzuelo, namely from Kamieneski et al. (2023, constrained by pairs of multiple images with JWST imaging data) and Caminha et al. (2023, cluster-scale model of El Gordo using HST and MUSE data sets). Caustics from our models are shown with dashed lines for clarity. The left panel is focused on the image plane, while the right panel is focused on the source region predicted by models including the cluster mass distribution. Note how the position of the caustics for models that do not include the cluster contribution (i.e., the Kamieneski et al. (2023) and “Shear only” models) is aligned with the lens (left panel), compared to the caustics of other models (right panel). The predicted (unlensed) surface brightness of the lens and source galaxies from the best-fit “Cluster+Shear” model (red curves) is shown in the background. North and East directions are indicated by the arrows.
@@
Although using pairs of multiply imaged features as constraints, the lens model of Kamieneski et al. (2023) is qualitatively very similar to ours. We show their critical lines and caustics [^11] An offset of $+0\aas@@fstack{\prime\prime}25\ {\rm E},+0\aas@@fstack{\prime\prime}01\ {\rm N}$ have been added to the critical lines and caustics of Kamieneski et al. (2023) for proper comparison with our models (P. Kamieneski, private communication). We have also found a similar offset between deflectors positions reported in Caminha et al. (2023) and our fitted light profile positions. in comparison to ours in Fig. 8. As expected, the main differences in the critical lines are directly related to the amount and location of the multiple images used in Kamieneski et al. (2023). A striking example is the westernmost part of the arc—around position $(2\aas@@fstack{\prime\prime}2,1^{\prime\prime})$—, for which Kamieneski et al. (2023) used a pair of multiple images which are probably unresolved by JWST and appears as point sources (images 2b and 2c in Fig. 9). In between these two images, all critical lines are overlapping, showing that both the arc and point-like constraints locally give similar results, despite our extended source model not including any point source components. On the contrary, in locations where either pairs of images are missing (e.g., south-western part of the arc) or inaccurate multiple images are used (northern part of the arc), clear differences do arise. In particular, the zoom-in inset in Fig. 8 focuses on a region with many multiple images of central regions of the source galaxy and shows deviations between the critical lines from different models. As we discuss in more detail in Sects. 5.4 and 5.5, we find that one feature in the arc may have been incorrectly attributed to a lensed image of the bright source bulge. We believe that this incorrect attribution causes the deviation between critical lines in the top right corner of the inset in Fig. 9, despite a region with a large number of lensing constraints. Focusing now on the caustics, we see a clear offset in the main astroid caustic between the model of Kamieneski et al. (2023) and our “Shear only” model. This result may be surprising at first, since both models are similarly parametrized (power-law profiles embedded in an external shear). However, this offset could be explained by the fact that our models including external shear all predict a non-zero mass for L3, as well as differences in the lensing constraints between the two approaches.
@@
We also compare in Fig. 8 our critical curves and caustics to the ones predicted by the cluster model. The cluster model was constructed without any constraints from El Anzuelo hence L1, L2 and L3 are mainly constrained through scaling relations based on their photometry (Caminha et al. 2023). We find that critical lines between our galaxy-scale model and this cluster-scale overall agree well, in particular in the vicinity L1 and L2. As shown in the inset of Fig. 8, the characteristic feature in the critical lines is comparable, meaning that qualitatively, the relative masses of L1 and L2 are consistent among the models. The cluster model displays rounder critical lines than our model, as expected from the use of spherically symmetric mass profiles. The discrepancy is the largest in the vicinity of L3, as this galaxy lies further away from the arcs. The cluster mass model assigns a larger mass to L3, which is also visible in source plane, where the eastern astroid caustic is significantly larger than in our models. While investigating this discrepancy is beyond the scope of this work, we simply mention that it may relate to the specific morphology of L3, which appears more elongated than L1 and L2 in the NIRCam data (see also the light models in Fig. 14). In addition, we notice that the light profile of L3 is very steep, as it can be seen from the prominent PSF spikes. Such a surface brightness distribution can be the sign of a disk component superimposed to a bright and compact bulge. In this case, a single-component mass profile like the SIE or dPIE might not be sufficiently accurate (for disk structures in lensing galaxies, see e.g., Hsueh et al. 2018).
Figure 9: Multiple images of peculiar features in the background source galaxy. Top row: families of multiple images in image plane together with the critical lines and the observed flux from the arc (after lens subtraction in filters F [^4] W ad F200W). Also shown are the positions of the three deflectors with pink crosses (residuals remaining after lens light subtraction are visible around L1 and L2 positions), as well as the critical lines with dashed white lines. Our model also predicts images at the position of L1 and L2, although there are not shown here to limit clutter. Bottom row: corresponding image positions on the source plane, superimposed onto one sample of our source model. The tangential caustic is shown with dotted white lines. Note that each column has its own dynamic range. Scale bars indicate the angular (in arcsec) and physical (in kpc) scales in image and source plane (assuming $z_{\rm s}=2.291$), respectively.
@@ 5.3 Constraints on mass model parameters
In this section we present our quantitative constraints on the mass distribution of the main deflectors. We show in Fig. 10 the joint posterior distributions of a subset of key mass model parameters, for each of the three model families (see Appendix C.2 for the full corner plot). We also summarize in Table 2 all constraints on mass model parameters for the fiducial “Cluster+shear” model. As expected, the Einstein radius of L1 and L2 as well the external shear strength significantly decrease after including the cluster mass distribution. Unlike the model of Kamieneski et al. (2023) composed of SIE profiles embedded in an external shear (which is the most similar to our “Shear only” model) the mass of L3 is constrained by the imaging data, except for the “Cluster only” model family. Our fiducial model predicts $\theta_{\rm E,L3}=0\aas@@fstack{\prime\prime}32\pm 0\aas@@fstack{\prime\prime}03$, which is $\sim 3\sigma$ away from the stellar kinematics estimate $\theta_{\rm E,SIS}=0\aas@@fstack{\prime\prime}49\pm 0\aas@@fstack{\prime\prime}05$. Therefore, the prior from stellar kinematics appears rather unconstraining compared to the imaging data, despite the angular distance between L3 and the lensing features.
@@
In Appendix C.3 we extensively check the robustness of our constraints against different modeling choices. These checks include the choice of VI sampler, the effect of inaccurate lens light model and the uncertainty associated to the cluster-scale model. In particular, we note that our models very weakly dependent on the source resolution, both in terms of constraints on mass model parameters and morphology of the reconstructed source. We believe this low dependence on the source resolution is a direct consequence of the efficient regularization of our correlated field source model, which allows us to reach high-resolution without compromising on computation time and model quality.
Finally, we find our lens model to be sensitive to changes in the cluster-scale mass distribution, as drawing different posterior samples from the cluster model of Caminha et al. (2023) lead to measurable changes in $\chi^{2}_{\nu}$. In particular, some posterior samples from the cluster model lead to a lower $\chi^{2}_{\nu}$ than with the best-fit cluster model we use in our fiducial model. Therefore, El Anzuelo lensing features may be used to discriminate between different cluster models, thus providing additional constraints on the mass distribution of large-scale components of El Gordo (i.e., its dark matter halo). Although this particular result demands further investigation, it shows that extended arcs, although primarily constraining the structure of individual cluster members, may also provide useful constraints on the mass distribution at the scale of the cluster.